# Reference-based Metrics Disprove Themselves in Question Generation

### v1

@@ 1 Introduction
Figure 1: Normalized value of different evaluation metrics for four types of candidate questions against the same reference (RefQ) in the HotpotQA dataset Yang et al. (2018). A good metric should give a significantly higher score to Group 1 than the other three groups. It can be seen that current reference-based metrics (in colors except red and blue) can identify the random questions (Group 4). However, they have trouble separating questions from Group 1 and Group 3. RQUGE, although capable of identifying two groups of questions that violate naturalness (Group 3) and answerability (Group 4), assigns a higher score for Group 2, which lacks complexity, than for Group 1. Our proposed metric (in red), prioritizing essential criteria of a question, can effectively distinguish all four groups of candidates while maintaining the highest rating for the valid questions.
@@
To evaluate QG performance, reference-based metrics are widely used, which assess a machine-generated question against a human-written reference. The metrics are calculated either at the word level such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005), or in the embedding space such as BERTScore Zhang et al. (2019). The challenges of using these evaluation metrics speak to the metrics themselves, considering word overlaps and/or semantic similarity between the generated question and the reference. In this sense, a QG model can “cheat” on the metrics by using many similar words to the reference, but ignoring essential components of a question. Mohammadshahi et al. questioned the effectiveness of the reference-based metrics, developed a QA model, and defined a new metric named “answerability” or RQUGE. Though they showed a slightly higher correlation with human preference, the failure of reference-based metrics was not studied, and the new metric’s effectiveness is sensitive to the QA model’s training and limited to its ability.
@@
To disprove existing metrics, the challenge can be traced to the lack of diverse references for benchmark datasets. Previous works have shown that with access to a more diverse pool of references, the problem of poor correlation for these metrics can be mitigated Freitag et al. (2020); Tang et al. (2023). However, QG benchmarks often contain one human-written ground-truth per example.
@@
Our study starts from collecting another set of human-written references for two QG benchmarks, following their standard annotation instructions. We also collect three groups of candidate questions, each lacking in an essential aspect of a question, for comparison. We study how five reference-based metrics, namely BLEU-4 Papineni et al. (2002), BLEURT Sellam et al. (2020), ROUGE-L Lin (2004), BERTScore Zhang et al. (2019), and Q-BLEU Nema and Khapra (2018), and two reference-free metrics, QAScore Ji et al. (2022), and RQUGE Mohammadshahi et al. (2023), score the four groups of questions. Figure 1 highlights the incompetency of current QG metrics in distinguishing the new reference (a valid question; see Group 1) from a less-complex-than-referenced question (Group 2), a non-question sentence that uses similar words (Group 3) or a randomly-selected question from training set (Group 4). Although these metrics tend to give higher scores for the new references than random questions, it remains challenging to separate them from the other less desirable candidates.
@@
Based on the above observations, we assert the failure of reference-based metrics in QG evaluation. We propose a shift to an evaluation mechanism that addresses essential criteria of a question that current metrics do not: (1) Naturalness: how natural the question sound Wang et al. (2020); Bi et al. (2021), (2) Answerability: whether the question is grounded to the given answer Ushio et al. (2022); Ji et al. (2022); Nema and Khapra (2018); Mohammadshahi et al. (2023), and (3) Complexity: how likely it requires inferencing and synthesizing information Wang et al. (2020); Bi et al. (2021). These criteria are not constrained to the syntactic and semantic structure of a single reference question. Thus, they can help tackle the challenges of evaluating question quality without access to a diverse set of references.
[…]
@@
We name the three-dimensional metric NACo. The red bars (the leftmost in each group) in Figure 1 show that NACo successfully distinguishes the valid questions (i.e., new human-written reference) from the other three groups with significant margins. Reference-based metrics are so heavily influenced by the presence of overlapping words between the original reference and an invalid candidate that they even prefer the invalid candidate that NACo assigns a significantly lower score.
@@
  • We produce an additional set of human-written questions to current QG benchmarks, and show the unreliability of reference-based metrics in reflecting the quality of generated questions.
@@ 2 Failure of Reference-based QG Metrics
2.1 Data Collection
In this study, we question the reliability of current evaluation metrics for question generation. To do so, we collect a new set of valid human-written questions and three other sets of questions for benchmark datasets, each violating an important criterion of a question addressed by NACo. Figure 1 illustrates our data collection process.
@@ 2.1 Data Collection
We perform the additional annotation on two datasets: (1) 748 test examples of SQuAD Rajpurkar et al. (2016), and (2) 48 test examples of HotpotQA Yang et al. (2018). For the HotpotQA sample, we also collected three other sets of questions, each violating an aspect required by the reference questions.
@@
Group 2: Single-hop questions for a multi-hop QG benchmark: This group of candidate questions targets the multi-hop characteristic of HotpotQA where the ground-truth questions are formed based on two passages. Specifically, we selected one from the original two passages that contains the answer span. We then asked GPT-3.5 to generate a question based on this single context passage. Next, we reviewed the questions for grammar, clarity, relevance to the passage, independence from external knowledge, and a logical path to the answer. We hypothesize that current QA-based metrics such as RQUGE Mohammadshahi et al. (2023) and QAScore Ji et al. (2022) fail to prefer the two-hop human-written questions to this set of single-hop questions.
@@
Group 3: Non-questions that use the same words as the reference: For this group of questions, we asked GPT-3.5 to generate a sentence based on the passages and use as many words from the same list of clues given to our annotators. We added a constraint such that the generated sentence cannot be in the form of a question. We then manually went through the generated sentences to ensure that no hallucinations were in place. In this sense, we produced a group of candidates that does not satisfy the most basic linguistic requirement of a question, naturalness, but still manages to contain many similar words as the ground-truth questions. We hypothesize that the majority of current evaluation metrics, especially those that rely mainly on word overlaps, fail to distinguish between the valid human-written questions in Group 1 and this group of non-questions.
@@
Group 4: Random questions from the training set: The final set of candidate questions comes randomly from the training set of the benchmark. In the example illustrated in Figure 1, the random question cannot be grounded to the given answer The Seduction of Hillary Rodham. Specifically, the answer to this candidate question is Robin McLaurin Williams, which is completely irrelevant to the given answer. In this sense, this group of candidate questions violates the answerability aspect of an ideal candidate. We hypothesize that the majority of current evaluation metrics can effectively recognize this set of random questions.
2.2 Study Design and Results
Previous studies aiming to challenge the effectiveness of reference-based metrics in QG typically rely on human evaluation. That is, they investigate whether the scores given to generated questions by QG metrics are highly correlated with the scores given by human evaluators Mohammadshahi et al. (2023); Ji et al. (2022). Unlike these studies, our research adopts a different approach during the data collection phase for benchmark datasets. Specifically, we replicate the data collection procedure of SQuAD and HotpotQA to collect new references, referred to as Group 1. Our focus is on determining if the newly collected references, when evaluated as candidates against the original references, receive high ratings from existing reference-based metrics. For the HotpotQA dataset, we extended our collection to include three additional groups of candidate questions considered less desirable (groups 2, 3, and 4) to ensure comprehensive comparisons. An effective and robust evaluation metric should assign a significantly higher score for questions in Group 1 compared to those in other groups.
@@ 2.2 Study Design and Results
Figure 1 shows the average normalized scores given by reference-based metrics to the four groups of candidate questions, all based on the same references. We find that all reference-based metrics, BLEU, ROUGE-L, BLEURT, Q-BLEU, and BERTScore, can effectively distinguish Group 4 (random questions) from the other groups, assigning it significantly lower score. For instance, the average ROUGE-L score for Group 4 is $0.11$, compared to $0.31$ for Group 1, $0.25$ for Group 2, and $0.39$ for Group 3, with a minimum difference of $14\%$ from the scores of the other groups.
@@
Figure 1 also reveals issues with the reference-based metrics in accurately assessing Groups 1, 2, and 3. Notably, for all five reference-based metrics, Group 3, non-question sentences with wording similar to the references, receives the highest average score. For example, the ROUGE-L metric scores a non-question sentence that uses similar wording to the reference on average $8\%$ higher than a new reference produced by our annotators, and $14\%$ higher than a perfectly answerable question requiring less reasoning than the reference. This indicates a flaw in reference-based metrics, as candidates that do not form coherent questions should not receive higher scores than those that do.
@@ 3 NACo: A Novel Multi-dimensional Reference-free QG Metric
We propose NACo, which leverages prompting and Chain-of-Thought (CoT) reasoning Wei et al. (2022) in answering the questions to be evaluated to obtain a score for each of the following criteria: Naturalness, Answerability, and Complexity. More specifically, given the relevant context passage(s) and a question, we instruct an LLM to answer the question based on the passage(s) as follows:
@@
If a candidate question passes the initial naturalness and answerability evaluation, we determine whether its complexity aligns with expected standards. Importantly, complexity expectations vary among domains and datasets. For instance, in the HotpotQA dataset, questions that require multi-hop reasoning might be preferred over simpler, single-hop questions. This observation may not apply for other datasets. In this sense, NACo relies on a subset of examples from the specific dataset to find the expected complexity of a question in that dataset. Specifically, we perform the above CoT-QA process to obtain the complexity of the references. Expected complexity is then defined by the most common number of reasoning steps needed by the LLM to answer a reference question. Our experiments using 750 examples from the training set find that for GPT-3.5, the expected complexity is $2$ for questions in SQuAD, and $3$ for questions in HotpotQA. Subsequently, NACo measures the similarity, denoted by $c_{cand}$, between the complexity of the candidate question and the expected complexity. Overall, NACo is a weighted combination of $n_{cand}$, $a_{cand}$, and $c_{cand}$. In our experiments, we adopt a fair weight $\frac{1}{3}$ for each criterion.
@@ 4 Experiments
4.1 Experimental Setup
We compare the evaluation capacity of NACo with that of current QG metrics on four QG models and three sets of human-validated references.
Generative Language Models like BART Lewis et al. (2020) and T5 Raffel et al. (2020) are current state-of-the-art QG performers on reference-based metrics Ushio et al. (2022). We fine-tune BART-base using the training set, following the method introduced by Chan and Fan. We refer to this model as BART-base. We also produce another version of BART-base that explores the effectiveness of clue-guided QG Liu et al. (2020). Specifically, before finetuning, we highlight the ground-truth clues used by reference questions (RefQ) in the context given as input to BART-base. We refer to this version as BART-clue-RefQ. In addition, we use GPT-3.5 to generate questions for the test examples through zero-shot and few-shot prompting. In the few-shot setting, we randomly select 10 examples from the training set of the dataset as demonstrations. We refer to these models as GPT-3.5 (zero-shot) and GPT-3.5 (few-shot).
[…]
4.2 Results
QG Competitor Ref-based metrics Ref-free metrics Our metric
B B-RT R-L BSc Q-B QA-S R-Q Na. Ans. Cmp. NACo
@@ 4.2 Results
BART-base 19.53 -0.28 44.79 92.13 36.94 -0.37 4.62 99.60 66.75 69.15 73.30
GPT-3.5 (few-shot) 18.06 -0.23 43.58 92.18 36.48 -0.37 4.56 100.00 65.23 69.63 73.67
BART-clue-RefQ 31.91 0.07 59.92 94.37 52.33 -0.38 4.56 99.73 63.94 69.24 69.97
@@
RefQ 100.00 100.00 100.00 100.00 100.00 -0.38 4.89 99.87 68.67 68.93 75.09
AnnoQ 12.78 -0.31 37.83 91.52 31.32 -0.37 4.71 99.73 67.70 70.33 74.01
AnnoQ-clue-RefQ 27.43 0.04 53.62 93.85 46.89 -0.38 4.76 99.87 66.34 70.36 74.21
Table 1: SQuAD - Performance of different QG methods on NACo and other existing metrics. The evaluation uses original SQuAD questions (RefQ) as references, with GPT-3.5 as the underlying QA system. Baseline metrics include BLEU (B), BLEURT (B-RT), ROUGE-L (R-L), BERTScore (BSc), Q-BLEU (Q-B), QAScore (QA-S), and RQUGE (R-Q). We also report individual score for each criterion Naturalness (Na.), Answerability (Ans), and Complexity (Cmp.) as measured by our NACo metric. The notation introduced here is consistent across subsequent tables for clarity and ease of comparison. The highest and second-highest scores (not including references for reference-based metrics) are highlighted with bold and underline markers, respectively.
QG Competitor Ref-based metrics Ref-free metrics Our metric
B B-RT R-L BSc Q-B QA-S R-Q Na. Ans. Cmp. NACo
LM-generated
BART-base 12.97 -0.79 34.38 87.88 27.67 -0.28 2.78 64.58 31.00 68.56 41.25
GPT-3.5 (zero-shot) 7.94 -0.95 26.05 87.38 17.00 -0.28 4.25 100.00 71.55 70.72 73.60
GPT-3.5 (few-shot) 10.53 -0.87 27.51 87.76 18.67 -0.27 4.21 99.31 67.31 72.64 71.09
BART-clue-RefQ 31.90 -0.10 58.77 92.61 51.98 -0.28 3.31 90.28 46.84 79.41 63.17
Human-validated
RefQ 100.00 100.00 100.00 100.00 100.00 -0.27 4.21 99.31 59.62 83.02 74.33
AnnoQ 13.64 -0.70 30.94 89.22 26.95 -0.27 4.32 100.00 76.11 82.05 83.43
AnnoQ-clue-RefQ 35.59 -0.06 52.95 92.15 51.76 -0.27 4.44 100.00 71.46 81.46 81.60
Table 2: HotpotQA - Performance of different QG methods on NACo and other existing metrics. The evaluation uses original HotpotQA questions (RefQ) as references, with GPT-3.5 as the underlying QA system. The highest and second-highest scores (not including references for reference-based metrics) are highlighted with bold and underline markers, respectively.
4.2.1 Failure of reference-based metrics
We report the performance of QG competitors on different evaluation metrics, including NACo. For reference-based metrics, we report their values using RefQ as the reference in Table 1 and 2. Values of reference-based metrics when AnnoQ are used as references are available in Appendix A.3.
Even though RefQ, AnnoQ, and AnnoQ-clue-RefQ are all qualified as valid questions, a large gap exists between them when reference-based metrics are in use. Specifically, in the SQuAD dataset, using RefQ as a reference, the BLEU scores for RefQ, AnnoQ, and AnnoQ-clue-RefQ are 100, 12.78, and 27.43, respectively, as shown in Table 1. Conversely, the gap in the reference-free metrics like RQUGE and NACo among these different sets of human-validated questions is much narrower when given the same input. Table 1 presents the NACo scores for RefQ, AnnoQ, and AnnoQ-clue-RefQ in the SQuAD dataset as 75.09, 74.01, and 74.21, respectively. Similar patterns are observed in the HotpotQA dataset, as detailed in Table 2.
According to reference-based metrics, models that learn from training data either through finetuning (like BART-base) or demonstration (like GPT-3.5) are scored significantly higher than our annotators, who lack access to the training data. For instance, in the case of SQuAD, BART-base is scored higher than AnnoQ by almost $7\%$ according to BLEU-4, reported in Table 1. As reference-based metrics measure syntactic and semantic similarity, the use of a single reference can disqualify our annotated questions from being considered reference materials, resulting in a misleading portrayal of a valid group of candidate questions.
4.2.2 Effectiveness of NACo
In this section, we demonstrate NACo’s superiority over both current reference-based and reference-free metrics. Referring to our analysis of four groups of candidate questions for HotpotQA, as illustrated in Figure 1, NACo uniquely succeeds in separating all four groups by significant margins, unlike the seven existing metrics. The newly collected multi-hop questions in Group 1, which meet all critical criteria for HotpotQA questions, achieve the highest average NACo score of $0.83$. They are followed by the questions in Group 2, lacking in complexity, with a score of $0.75$; Group 3, lacking in naturalness, with a score of $0.19$; and Group 4, lacking in answerability, with a score of $0.03$.
[…]
In the following section, we further validate our claims regarding NACo’s effectiveness in QG assessment through human evaluation.
4.3 Human Evaluation
@@
NACo 0.40 0.25 0.19 0.70 0.59 0.49 0.44 0.34 0.27 0.68 0.49 0.38
@@
Table 3: Correlation between human assessments and automated evaluation metrics as indicated by Pearson $r$, Spearman $\rho$, and Kendall $\tau$ correlation coefficients. The highest and second-highest scores are highlighted with bold and underline markers, respectively.

### v2

@@ 1 Introduction
Figure 1: Normalized value of different evaluation metrics for four types of candidate questions against the same reference (RefQ) in the HotpotQA dataset Yang et al. (2018). Ideally, metrics should score Group 1 highest. Current QG metrics, except for NACo (ours) and RQUGE, primarily recognize random questions (Group 4) but fail to differentiate between Groups 1 and 3 (note the red and green bars). RQUGE, successfully identifies groups violating naturalness (Group 3) and answerability (Group 4), assigns a higher score for Group 2, which lacks complexity, than for Group 1. Our metric, shown in the leftmost bar group, prioritizing essential criteria of a question, can effectively distinguish all four groups of candidates while maintaining the highest rating for the valid questions.
@@
To evaluate QG performance, reference-based metrics are widely used, which assess a machine-generated question against a human-written reference. The metrics are calculated either at the word level such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005), or in the embedding space such as BERTScore Zhang et al. (2019). The challenges of using these evaluation metrics speak to the metrics themselves, considering word overlaps and/or semantic similarity between the generated question and the reference. In this sense, a QG model can “cheat” on the metrics by using many similar words to the reference, but ignoring essential components of a question. Mohammadshahi et al. questioned the effectiveness of reference-based metrics, developed a QA model, and defined a new metric named “answerability” or RQUGE. Though they showed a higher correlation with human preference, the failure of reference-based metrics was not studied, and the new metric’s effectiveness is sensitive to the QA model’s training and limited to its ability.
@@
To disprove existing metrics, the challenge can be traced to the lack of diverse references for benchmark datasets. Previous works have shown that with access to a more diverse pool of references, the problem of poor correlation for these metrics can be mitigated Freitag et al. (2020); Oh et al. (2023); Tang et al. (2023). However, QG benchmarks often contain only one human-written ground-truth per example.
@@
Our study starts from collecting another set of human-written references for two QG benchmarks, following their standard annotation instructions. Besides the new references, we collect three groups of candidate questions, each lacking in an essential aspect of a question, for comparison. We study how five reference-based metrics, namely BLEU-4 Papineni et al. (2002), BLEURT Sellam et al. (2020), ROUGE-L Lin (2004), BERTScore Zhang et al. (2019), and Q-BLEU Nema and Khapra (2018), and two reference-free metrics, QAScore Ji et al. (2022), and RQUGE Mohammadshahi et al. (2023), score the four groups of questions. Fig. 1 highlights the incompetency of current QG metrics in distinguishing the new reference (a valid question; see Group 1) from a less-complex-than-referenced question (Group 2), a non-question sentence that uses similar words (Group 3) or a randomly-selected question from training set (Group 4). Although these metrics tend to give higher scores for the new references than random questions, it remains challenging to separate them from the other less desirable candidates.
@@
Based on the above observations, we assert the failure of reference-based metrics in QG evaluation. We propose a shift to an evaluation mechanism that addresses essential criteria of a question that current metrics neglect: (1) Naturalness: how natural the question sounds Wang et al. (2020); Bi et al. (2021), (2) Answerability: whether the question is grounded to the given answer Ushio et al. (2022); Ji et al. (2022); Nema and Khapra (2018); Mohammadshahi et al. (2023), and (3) Complexity: how likely it requires inferencing and synthesizing information Wang et al. (2020); Bi et al. (2021). These criteria are not constrained to the syntactic and semantic structure of a single reference question. Thus, they address the challenges of evaluating question quality without access to a diverse set of references.
To overcome the limitation of the answerability measure in RQUGE and implement the other two measures, we utilize large language models (LLMs), which have demonstrated potential utility in data annotation tasks Liu et al. (2023); Wang et al. (2023); Lin and Chen (2023); Chiang and Lee (2023), and their Chain-of-Thought (CoT) Wei et al. (2022) process. We design CoT prompts for the LLM to directly measure the three criteria, as described in detail in §$\S$3.
@@
We name the three-dimensional metric NACo. The leftmost group of bars in Fig. 1 shows that NACo successfully distinguishes the valid questions (i.e., new human-written reference) from the other three groups with significant margins. Reference-based metrics are so heavily influenced by the presence of overlapping words between the original reference and an invalid candidate that they even prefer the invalid candidate that NACo assigns a significantly lower score.
@@
  • We produce an additional set of human-written questions to current QG benchmarks, and show the unreliability of reference-based metrics in reflecting question quality.
@@ 2 Failure of Reference-based QG Metrics
2.1 Study Design & Data Collection
Previous studies questionting the effectiveness of reference-based metrics in QG typically rely on human evaluation. That is, they investigate whether the scores given to generated questions by QG metrics are highly correlated with the scores given by human evaluators Mohammadshahi et al. (2023); Ji et al. (2022). Unlike these studies, our research adopts a different approach during the data collection phase for QG datasets. Specifically, we replicate the data collection procedure of the datasets to collect new references, referred to as Group 1. Our focus is on determining if the newly collected references, when evaluated as candidates against the original references, receive high ratings from existing metrics. In addition, we extended our collection procedure to include three additional groups of candidate questions considered less desirable (Groups 2, 3, and 4) to ensure comprehensive comparisons. An effective and robust evaluation metric should assign a significantly higher score for questions in Group 1 compared to those in other groups. Fig. 1 illustrates our data collection process.
@@ 2.1 Study Design & Data Collection
We perform the additional annotation on two popular QG benchmarks: (1) 748 test examples of SQuAD Rajpurkar et al. (2016), and (2) 48 test examples of HotpotQA Yang et al. (2018). To illustrate the application of our study, we collect another QG dataset in the educational domain, specifically from the TED-Ed learning platform [^1] https://ed.ted.com/. We further annotate 43 questions from this new dataset. More details about data collection and annotation for TedEdQA are provided in Appx. A.5.
@@
For the HotpotQA sample, we also collected three other sets of questions, each violating an aspect required by the reference questions.
@@
Group 2: Single-hop questions for a multi-hop QG benchmark: This group of candidate questions targets the multi-hop characteristic of HotpotQA where the ground-truth questions are formed based on two passages. Specifically, we selected one from the original two passages that contains the answer span. We then asked GPT-3.5 to generate a question based on this single passage. We reviewed the questions for grammar, clarity, relevance to the passage, independence from external knowledge, and a logical path to the answer.
@@
Group 3: Non-questions that use the same words as the reference: For this group of questions, we asked GPT-3.5 to generate a sentence based on the passages and use as many words from the same list of clues given to our annotators. We added a constraint such that the generated sentence cannot be in the form of a question. We then manually went through the generated sentences to ensure that no hallucinations were in place. In this sense, we produced a group of candidates that does not satisfy the most basic linguistic requirement of a question, naturalness, but still manages to contain many similar words as the ground-truth questions.
@@
Group 4: Random questions from the training set: The final set of candidate questions comes randomly from the training set of the benchmark. In the example illustrated in Fig. 1, the answer to this candidate question is Robin McLaurin Williams, which is completely irrelevant to the given answer The Seduction of Hillary Rodham. In this sense, this group of candidate questions violates the answerability aspect of an ideal candidate.
2.2 Results
@@
Fig. 1 shows the average normalized scores given by reference-based metrics to the four groups of candidate questions, all based on the same references. We find that all reference-based metrics, BLEU, ROUGE-L, BLEURT, Q-BLEU, and BERTScore, can effectively distinguish Group 4 (random questions) from the other groups, assigning it significantly lower score. For instance, the average ROUGE-L score for Group 4 is $0.11$, compared to $0.31$ for Group 1, $0.25$ for Group 2, and $0.39$ for Group 3, with a minimum difference of $14\%$ from the scores of the other groups.
@@
Fig. 1 also reveals issues with the reference-based metrics in accurately assessing Groups 1, 2, and 3. Notably, for all five reference-based metrics, Group 3, non-question sentences with wording similar to the references, receives the highest average score. For example, the ROUGE-L metric scores a non-question sentence that uses similar wording to the reference (green bar) on average $8\%$ higher than a new reference produced by our annotators (red bar), and $14\%$ higher than a perfectly answerable question requiring less reasoning than the reference (blue bar). This observation indicates a flaw in reference-based metrics, as candidates that do not form coherent questions should not receive higher scores than those that do.
@@ 3 NACo: A Novel Multi-dimensional Reference-free QG Metric
Based on extensive review of the human evaluation procedure in QG literature, detailed in Appx. A.1, we identify three essential criteria of a question: Naturalness, Answerability, and Complexity. We propose NACo, which leverages prompting and Chain-of-Thought (CoT) reasoning Wei et al. (2022) to obtain a score for each criterion. Specifically, given the relevant context passage(s) and a question, we instruct an LLM as follows:
@@
If a candidate question passes the initial naturalness and answerability evaluation, we determine whether its complexity aligns with expected standards for the domain and dataset. For example, in the HotpotQA dataset, questions that require multi-hop reasoning might be preferred over simpler, single-hop questions. This preference may not hold in other datasets. In this sense, NACo relies on a subset of examples from the specific dataset to find the expected complexity of a question in that dataset. Specifically, we perform the above CoT-QA process to obtain the complexity of the references. Expected complexity is then defined by the most common number of reasoning steps needed by the LLM to answer a reference question. Our experiments using 750 examples from the training set find that for GPT-3.5, the expected complexity is $2$ for questions in SQuAD, and $3$ for questions in HotpotQA. Subsequently, NACo measures the similarity, denoted by $c_{cand}$, between the complexity of the candidate question and the expected complexity.
Overall, NACo is a weighted combination of $n_{cand}$, $a_{cand}$, and $c_{cand}$. In our experiments, we adopt a fair weight $\frac{1}{3}$ for each criterion. We provide additional details on how $c_{cand}$ is computed and integrated into the final score in Appx. A.3
@@ 4 Experiments
QG Competitor Ref-based metrics NACo
B B-RT R-L BSc Q-B
@@
BART-base 19.53 -0.28 44.79 92.13 36.94 73.30
GPT-3.5 (few-shot) 18.06 -0.23 43.58 92.18 36.48 73.67
BART-clue-RefQ 31.91 0.07 59.92 94.37 52.33 69.97
@@
RefQ 100.00 1.00 100.00 100.00 100.00 75.09
AnnoQ 12.78 -0.31 37.83 91.52 31.32 74.01
AnnoQ-clue-RefQ 27.43 0.04 53.62 93.85 46.89 74.21
Table 1: SQuAD - Performance of different QG methods on NACo and other existing metrics. The evaluation uses original SQuAD questions (RefQ) as references, with GPT-3.5 as the underlying LLM. The highest and second-highest scores (not including references for reference-based metrics) are highlighted with bold and underline markers, respectively.
@@
NACo (Ours)
       Llama3-8B 0.43 0.33 0.27 0.60 0.51 0.43 0.44 0.36 0.29 0.65 0.49 0.38
       Mixtral-8x7B 0.37 0.29 0.23 0.58 0.52 0.43 0.41 0.27 0.21 0.59 0.43 0.33
       Claude-Haiku 0.47 0.26 0.21 0.67 0.49 0.40 0.53 0.44 0.36 0.73 0.54 0.43
       GPT3.5 0.40 0.25 0.19 0.70 0.59 0.49 0.44 0.34 0.27 0.68 0.49 0.38
       GPT4 0.55 0.29 0.25 0.72 0.56 0.48 0.56 0.38 0.36 0.8 0.53 0.44
@@
Table 2: Correlation between human assessments and automated evaluation metrics as indicated by Pearson $r$, Spearman $\rho$, and Kendall $\tau$ correlation coefficients. For reference-based metrics, we report the metric’s correlation with human judgment both when only the original reference is used and when adding the diversified references (w/ DivRef). The highest and second-highest scores are highlighted with bold and underline markers, respectively. Shaded regions indicate an improvement compared to the current state-of-the-art metric for that respective column.
4.1 Experimental Setup
Question generation competitors: We compare the evaluation capacity of NACo with that of current QG metrics on four QG models and three sets of human-validated references. Generative Language Models like BART Lewis et al. (2020) and T5 Raffel et al. (2020) are current state-of-the-art QG performers on reference-based metrics Ushio et al. (2022). We fine-tune BART-base using the training set, following the method introduced by Chan and Fan. We also produce another version of BART-base, BART-clue-RefQ, which highlight the ground-truth clues used by reference questions (RefQ) in the context given as input to BART-base (detailed in A.6). In addition, we use GPT-3.5 to generate questions for the test examples through zero-shot, and few-shot prompting. In the few-shot setting, we randomly select 10 examples from the training set of the dataset as demonstrations. Alongside the original reference questions provided by the datasets (RefQ), we use the annotated data detailed in §$\S$2 to obtain two human-validated competitors: AnnoQ, which contains the questions written by our annotators before given gold clues, and AnnoQ-clue-RefQ, which contains the gold-clue-guided questions written by our annotators.
Baselines: We compare the evaluation capacity of NACo with five reference-based metrics, including BLEU-4 (B) Papineni et al. (2002), BLEURT (B-RT) Sellam et al. (2020), ROUGE-L (R-L) Lin (2004), BERTScore (BSc) Zhang et al. (2019), and Q-BLEU (Q-B) Nema and Khapra (2018), and two reference-free metrics, QAScore (QA-S) Ji et al. (2022), and RQUGE (R-Q)Mohammadshahi et al. (2023). Tang et al. propose using LLM to diversify the limited references in benchmarks, demonstrating an improvement in the correlation between reference-based metrics and human judgment. We replicate this approach and report the evaluation performance of the five reference-based metrics both when only the original reference is used and when adding the diversified references.
NACo implementation: We provide the CoT prompt used in our experiments in Appx. A.2. We experimented with five underlying LLMs: Llama3-8B, Mixtral-8x7B, Claude3-Haiku, GPT3.5-turbo, and GPT4o.